# Scale-Aware Pre-Training for Human-Centric Visual Perception: Enabling Lightweight and Generalizable Models

### v1

@@ 4.3 Main Results
Table 3: Quantitative comparison with state-of-the-art pretraining methods on 3 multi-person visual understanding tasks. $\downarrow$ means the smaller value the better performance.
In this section, we compare SAIP with existing SSP methods across a wide range of HVP tasks. For comprehensive evaluations, we investigate three lightweight architectures, i.e. ViT_Tiny [8], StarNet_s3 [31] and TinyVit_5M [39].
Single-person Discrimination. As indicated in Table 1, SAIP achieves substantial performance improvement over existing methods. Using the ViT_Tiny as the backbone, SAIP outperforms previous best method SOLIDER by 2.0% and 6.4% on Market and MSMT17, 3.7% and 5.4% on CUHK-PEDES and ICFG-PEDES, 2.1% and 2.0% on PA100K and PETAzs, respectively. Furthermore, using a pretrained large model (i.e. PATH [35]) as the expert encoder, the performance improvements become much higher, achieving 3%-13%. This strong downstream performance indicates that the proposed SAIP serves not only as a SSP method but also an effective knowledge distillation solution. Similar improvements are also observed when utilizing StarNet or TinyVit as the backbone.
Single-person Dense Predictions. As illustrated in Table 2, we fine-tune the parameters of SAIP models across three dense predictions tasks. With the same parameter count and same model architecture, the SAIP models exhibit superior performance compared to the counterparts of state-of-the-art SSP methods. For example, the SAIP pretrained ViT_Tiny surpasses previous methods by 1.6%-3.9% in $AP$ scores for pose estimation task, by 2.1%-6.2% in $AP$ scores for landmark detection task, and by 6.3%-11.7% in $mIoU$ scores for part segmentation task. More interestingly, in the context of dense prediction tasks, employing a pretrained large model as the expert encoder does not yield a significant improvement, suggesting that visual patterns learned by SAIP are comparable to those learned from a pretrained large model.
Multi-person Understanding. Table 3 presents results for tasks that require to simultaneously solve person location, shape understanding and multi-granularity reasoning. Overall, SAIP surpasses previous SSP methods under the same setting. For example, SAIP family outperforms previous SSP methods by 1.0%-3.8% $AP$ scores for pedestrian detection, 2.3%-5.1% $mIoU$ scores for multiple human parsing, and 1.3%-6.0% $AP^{box}_{IoU+F_{1}}$ scores for part-level attribute parsing.
Qualitative Comparison. We seek to reveal the learned visual patterns by conducting a comparative visualization analysis across different SSP methods. Specifically, we examine two scenarios, i.e. single instance across two scales and multiple instances within a single image. The visual patterns learned by various SSP methods are presented in Fig. 4. Based on the visualization comparison, we have the following observations: First, most vision models, particularly those based on MIM (i.e., MAE and HAP), fail to capture visual patterns in these two scenarios. Second, SAIP-pretrained model demonstrates the ability to capture visual pattern for a person across multiple scales, but exhibit limitations in handling multiple instances. Third, when a stronger expert encoder is employed during pretraining, SAIP-pretrained model effectively captures meaningful patterns in both scenarios. These observations, combined with the quantitative comparisons, suggest a clear correlation, that is, vision models that capture more comprehensive human visual patterns consistently achieve superior downstream performance.
Based on these comparative experiments, one can conclude that proposed SAIP performs a general improvement across various HVP tasks and exhibits a strong generalization capability.
4.4 Ablation Study
In this section, we conduct ablation studies to investigate the major properties of SAIP. For fair comparison, we adopt the default settings, i.e. ViT_Tiny serving as the backbone model and LUP1M adopted as pretraining dataset.
@@ 4.4 Ablation Study
Table 4: Investigation of the effect of learning tasks in SAIP.
Effectiveness of the Devised Components in SAIP. The SAIP is designed with three learning objectives, i.e. cross-scale matching (CSM), cross-scale reconstruction (CSR) and cross-scale search (CSS). We conduct a component-wise analysis by progressively using each objective to pretrain lightweight models, seeking to reveal the impact of each learning objective. Randomly initialized ViT_Tiny (i.e. without pretraining) and its IN1K pretrained counterpart are used as baseline models. As reported in Table 4, the pretrained ViT_Tiny exhibits significantly superior performance compared to the randomly initialized variant, indicating the critical importance of pretraining. Furthermore, each cross-scale learning objective makes a positive contribution to performance, achieving results that are either superior or comparable to those of IN1K pretrained model. This indicates that a proper self-supervised learning strategy can yield better visual patterns than those derived from supervised learning. Additionally, it is evident that CSM, CSR, and CSS complement one another, collectively improving performance on downstream tasks. This illustrates that multi-task self-supervised learning is particularly beneficial for unlocking lightweight model’s potential to learn general human visual patterns.
Impact of the Pretraining Dataset. Existing works [2, 35, 23] emphasize the significance of the quantity of person images. In this work, we further extend the analysis by considering two additional dimensions of the pretraining dataset: the diversity of object categories and high fidelity of person images. To this end, we prepare five distinct datasets as listed in Table 5. In addition to ImageNet-1K (IN1K) [5], which contains 1000 object categories, we further utilize LUPerson dataset [10] to construct three datasets: LUP1M, LUP2M, and LUP4M with 1 million, 2 million, and 4 million person images, respectively. Notably, LUPerson contains significant noise. To develop a dataset with high fidelity person images, we collect 1 million images from 7 publicly available datasets (COCO, AIC, etc). For simplicity, the collected dataset is denoted as HP1M.
Comparison results in Table 5 indicate that improving fidelity of person images benefits model performance more effectively than increasing diversity, while merely increasing the quantity does not bring a noticeable improvement. Furthermore, combining LUP1M with IN1K and HP1M leads to the best performance, demonstrating that high fidelity person images with a good diversity are particularly important for pretraining lightweight models.
Training Efficiency in Downstream Tasks. To investigate the impact of SAIP on downstream fine-tuning, we fine-tune SAIP models using four distinct training schedules. As shown in Table 6, the optimal performance is attained using 50% of the fine-tuning epochs, presenting a rapid convergence. This indicates that SAIP is beneficial for accelerating convergence in fine-tuning stage, thereby establishing it as a computationally sustainable solution.
@@
Table 5: Investigating the impact of the data from three aspects: quantity, diversity and quality. Increasing the quantity does not bring a noticeable improvements.
@@
Table 6: Downstream finetuning. SAIP achieves optimal results with 50% finetuning epochs.
5 Conclusion
@@
In this paper, we identify two key limitations in adapting existing HVP models to real-world applications, i.e. specific visual patterns and excessively large model size. To handle these, we propose SAIP, the first self-supervised pretraining framework designed to establish lightweight and generalizable models for HVP. Extensive experiments on a wide range of HVP tasks demonstrate the effectiveness and strong generalization capabilities of SAIP. Furthermore, we highlight two useful findings related to pretraining lightweight vision models. These findings suggest that if a proper training objective is explored, lightweight architectures can serve as a good generalizable model for HVP. We hope this work can bring inspiration to the community.

### v2

@@ IV-B Implementation Details
Supervised Pretraining StarNet_S3 [30] 5.8M IN1k 89.2 45.3 50.9 49.4 36.9 33.8
ResNet18 [47] 11M IN1K 86.9 48.7 51.3 49.6 35.2 32.8
ViT_Tiny [32] 5.5M IN1K 86.9 50.0 47.5 46.1 37.5 34.8
EdgeNeXt [33] 5.6M IN1K 89.6 44.3 53.0 51.3 40.1 37.1
TinyViT_5M [34] 5.4M IN1K 88.9 44.6 51.7 50.4 38.5 35.3
Self-supervised Pretraining
@@
TABLE III: Quantitative comparison with state-of-the-art pretraining methods on 3 multi-person visual understanding tasks. $\downarrow$ means the smaller value the better performance.
@@
TABLE IV: Investigation of the effect of learning tasks.
@@
TABLE V: Investigating the impact of the data from three aspects: quantity, diversity and quality. Increasing the quantity does not bring a noticeable improvements.
@@
TABLE VI: Downstream finetuning. SAIP achieves optimal results with 50% finetuning epochs.
IV-C Main Results
In this section, we compare SAIP with existing SSP methods and supervised pretraining methods across a wide range of HVP tasks. For comprehensive evaluations, we investigate three lightweight architectures, i.e., ViT_Tiny [32], StarNet_s3 [30] and TinyVit_5M [34].
Single-person Discrimination. As indicated in Table I, SAIP achieves substantial performance improvement over existing methods. Using the ViT_Tiny as the backbone, SAIP outperforms previous best method SOLIDER by 2.0% and 6.4% on Market and MSMT17, 3.7% and 5.4% on CUHK-PEDES and ICFG-PEDES, 2.1% and 2.0% on PA100K and PETAzs, respectively. Notably, SAIP’s performance even surpasses that of supervised pretraining methods on these tasks. Furthermore, using a pretrained large model (i.e., PATH [10]) as the expert encoder, the performance improvements become much higher, achieving 3%-13%. This strong downstream performance indicates that the proposed SAIP serves not only as a SSP method but also an effective knowledge distillation solution. Similar improvements are also observed when utilizing StarNet or TinyVit as the backbone.
Single-person Dense Predictions. As illustrated in Table II, we fine-tune the parameters of SAIP models across three dense predictions tasks. With the same parameter count and same model architecture, the SAIP models exhibit superior performance compared to the counterparts of state-of-the-art SSP methods and supervised pretraining methods. For example, the SAIP pretrained ViT_Tiny surpasses previous methods by 1.6%-3.9% in $AP$ scores for pose estimation task, by 2.1%-6.2% in $AP$ scores for landmark detection task, and by 6.3%-11.7% in $mIoU$ scores for part segmentation task. More interestingly, in the context of dense prediction tasks, employing a pretrained large model as the expert encoder does not yield a significant improvement, suggesting that visual patterns learned by SAIP are comparable to those learned from a pretrained large model.
Multi-person Understanding. Table III presents results for tasks that require to simultaneously solve person location, shape understanding and multi-granularity reasoning. Overall, SAIP surpasses previous SSP methods and supervised pretraining methods under the same setting. For example, SAIP family outperforms previous SSP methods by 1.0%-3.8% $AP$ scores for pedestrian detection, 2.3%-5.1% $mIoU$ scores for multiple human parsing, and 1.3%-6.0% $AP^{box}_{IoU+F_{1}}$ scores for part-level attribute parsing.
Based on these comparative experiments, one can conclude that proposed SAIP performs a general improvement across various HVP tasks and exhibits strong generalization.
Figure 4: Training schedules. A longer training schedule does not give a noticeable improvement.
IV-D Ablation Study
In this section, we conduct ablation studies to investigate the major properties of SAIP. For fair comparison, we adopt the default settings, i.e., ViT_Tiny serving as the backbone model and LUP1M adopted as pretraining dataset.
Effectiveness of the Devised Components in SAIP. The SAIP is designed with three learning objectives, i.e., cross-scale matching (CSM), cross-scale reconstruction (CSR) and cross-scale search (CSS). We conduct a component-wise analysis by progressively using each objective to pretrain lightweight models, seeking to reveal the impact of each learning objective. Randomly initialized ViT_Tiny (i.e., without pretraining) and its IN1K pretrained counterpart are used as baseline models. As reported in Table IV, the pretrained ViT_Tiny exhibits significantly superior performance compared to the randomly initialized variant, indicating the critical importance of pretraining. Furthermore, each cross-scale learning objective makes a positive contribution to performance, achieving results that are either superior or comparable to those of IN1K pretrained model. This indicates that a proper self-supervised learning strategy can yield better visual patterns than those derived from supervised learning. Additionally, it is evident that CSM, CSR, and CSS complement one another, collectively improving performance on downstream tasks. This illustrates that multi-task self-supervised learning is particularly beneficial for unlocking lightweight model’s potential to learn general human visual patterns.
Impact of the Pretraining Dataset. Existing works [11, 10, 13] emphasize the significance of the quantity of person images. In this work, we further extend the analysis by considering two additional dimensions of the pretraining dataset: the diversity of object categories and high fidelity of person images. To this end, we prepare five distinct datasets as listed in Table V. In addition to ImageNet-1K (IN1K) [48], which contains 1000 object categories, we further utilize LUPerson dataset [16] to construct three datasets: LUP1M, LUP2M, and LUP4M with 1 million, 2 million, and 4 million person images, respectively. Notably, LUPerson contains significant noise. To develop a dataset with high fidelity person images, we collect 1 million images from 7 publicly available datasets (COCO, AIC, etc). For simplicity, the collected dataset is denoted as HP1M.
Comparison results in Table V indicate that improving fidelity of person images benefits model performance more effectively than increasing diversity, while merely increasing the quantity does not bring a noticeable improvement. Furthermore, combining LUP1M with IN1K and HP1M leads to the best performance, demonstrating that high fidelity person images with a good diversity are particularly important for pretraining lightweight models.
Training Efficiency in Downstream Tasks. To investigate the impact of SAIP on downstream fine-tuning, we fine-tune SAIP models using four distinct training schedules. As shown in Table VI, the optimal performance is attained using 50% of the fine-tuning epochs, presenting a rapid convergence. This indicates that SAIP is beneficial for accelerating convergence in fine-tuning stage, thereby establishing it as a computationally sustainable solution.
Impact of Pretraining Epochs. In general, large-scale pretraining often brings heavy training burden, which significantly limit affordability of SSP methods. However, we surprisingly find that proposed SAIP is beneficial for reducing training costs in pretraining stage. As shown in Fig 4, SAIP achieves optimal downstream performance at early stage of pretraining schedule, i.e., at 200-th epoch rather than 300-th epoch.
Method Humanart Chimpact-Pose AP-10K
DINO [1] 65.7 16.1 58.0
MAE [2] 65.1 13.9 48.8
MAE + DINO [5] 67.4 18.2 59.5
HAP [8] 66.0 13.2 50.5
SOLIDER [11] 66.7 16.0 57.4
SAIP (Ours) 67.5 18.2 60.8
TABLE VII: Investigating generalization capability of HVP models by testing them on unseen domains, including: (1) person images with unseen styles, (2) animal images, and (3) common object images.
Cross-domain Generalizability of SAIP. We examine cross-domain generalization capabilities using leave-one-domain-out evaluation, testing how self-supervised pretrained models transfer to novel domains absent from pretraining data. In our experimental design, we employ natural person images for pretraining and evaluate domain generalization on three distinct unseen domains: 1) stylized person images with various styles such as cartoons and sketches; 2) natural animal images, and 3) natural common object images. We conduct this investigation using three representative datasets, i.e., Humanart [49], Chimpact-Pose [50] and AP-10K [51]. Testing results presented in Tab VII indicate that our approach surpasses prior SSP methods. This demonstrates that the SAIP pretrained model learns general representations that are robust to domain changes and exhibits strong cross-domain generalization capability.
Qualitative Comparison. We seek to reveal the learned visual patterns by conducting a comparative visualization analysis across different SSP methods. Specifically, we examine two scenarios: single instance across two scales and multiple instances within a single image. The visual patterns learned by various SSP methods are presented in Fig. 5. Based on the visualization comparison, we have the following observations: First, most vision models, particularly those based on MIM (i.e., MAE and HAP), fail to capture visual patterns in both two scenarios. Second, SAIP-pretrained model demonstrates the ability to capture visual pattern for a person across multiple scales, but exhibit limitations in handling multiple instances. Third, when pretrained with a stronger expert encoder, SAIP-pretrained model effectively captures meaningful patterns in both scenarios. These observations, corroborated by quantitative comparisons, reveal a clear correlation: vision models that capture more diverse human visual patterns consistently achieve superior downstream performance.
V Conclusion
@@
In this paper, we identify two key limitations in adapting existing HVP models to real-world applications, i.e., specific visual patterns and excessively large model size. To handle these, we propose SAIP, the first self-supervised pretraining framework designed to establish lightweight and generalizable models for HVP. Extensive experiments on a wide range of HVP tasks demonstrate the effectiveness and strong generalization capabilities of SAIP. Furthermore, we highlight two useful findings related to pretraining lightweight vision models. These findings suggest that if a proper training objective is explored, lightweight architectures can serve as a good generalizable model for HVP. We hope this work can bring inspiration to the community.
A. SAIP Pretraining Details
We first present the implementation details of major components of SAIP, including the architectures of tasks decoders, pretraining settings and pretraining datasets.
CSM decoder consists of a Multi-Layer Perceptron (MLP) block, a $l_{2}$ normalization layer and a projection layer, following the design principles of DINO [1]. The MLP block, which comprises three non-linear layers, transforms image tokens of $D$ dimensions (i.e., 192) into a latent representation of 256 dimensions. Subsequently, the latent representation is normalized by $l_{2}$ normalization layer and further mapped to a high-dimensional instance representation (i.e., 65536) by the projection layer.
CSR decoder is a stack of 6 transformer blocks, each incorporating a multi-head attention block, LayerNorm layers applied before and after attention block, and nonlinear layers. The CSR decoder processes sequential tokens encoded by the image encoder, including tokens from the anchor image and visible tokens from the masked image.
CSS decoder includes two attention blocks identical to those in the CSR decoder. In addition, a cross-attention block, four convolution layers and one predictive linear layer are appended to the attention blocks. Specifically, the convolution layers and the predictive linear layer are used for binary mask extraction. The binary mask represents a rectangular region corresponding to an instance.
Pretraining setting. The default pretraining settings are presented in Table  VIII. All transformer blocks are initialized using the Xavier uniform. The learning rate is linearly warmed up over the first 10 epochs and subsequently decayed using a cosine schedule. The anchor image size is set to $256\times 128$, while the multi-person image has a size of $224\times 224$.
Config Value
Optimizer AdamW
Base Learning Rate 2.5e-4
Weight Decay 0.05
Optimizer Momentum $\beta_{1},\beta_{2}=0.9,0.95$
Batch Size per gpu 256
Learning Rate Schedule Cosine scheduler
Warmup Epochs 10
TABLE VIII: Pretraining setting.
Pretraining dataset. In the paper, we investigate the influence of pretraining data by constructing 5 different person-centric datasets, i.e., IN1K, LUP1M, LUP2M, LUP4M and HP1M. Specifically, LUPxM is derived from LUPerson dataset [16], and HP1M is a high-quality dataset collected from 7 publicly available datasets, i.e., AIC [52], COCO [42], CIHP [26], MHPv2 [53], CrowdHuman [44], Fashionpedia [27], and HumanArt [49]. Notably, these datasets provide manually annotated bounding boxes. Therefore, we collect images from the training sets of these 7 datasets, and use the provided bounding boxes to crop person instances from collected images, resulting in approximately 1 million high-fidelity person images. Detailed statistics of HP1M are presented in Table IX.
dataset #images
AIC [52] 378352
COCO [42] 149813
CIHP [26] 93207
MHPv2 [53] 41056
CrowdHuman [44] 428109
Fashionpedia [27] 45623
HumanArt [49] 81976
Total 1218136
TABLE IX: The statistics of HP1M.
B. Downstream Fine-tuning Details
For the evaluation of pretrained methods, we adopt representative solutions in downstream tasks as baselines, replacing their backbones with the pretrained models.
I2I person ReID. We use the TransReID [18] implemented in [11] as the basic solution for I2I ReID. All models are fine-tuned on Market1501 [36] train and MSMT17 [37] train respectively, and further evaluated on Market1501 test and MSMT17 test. The training process consists of 120 epochs with a default learning rate of 2e-4. During both training phase and testing phase, the image size is set to $256\times 128$.
T2I person ReID. We use the IRRA [19] with its publicly available code base to evaluate pretrained models. All pretrained models are fine-tuned on CUHK-PEDES [38] train or ICFG-PEDES [39] train, and tested on CUHK-PEDES test or ICFG-PEDES test. Each model is fine-tuned for 60 epochs, with the image size during both training phase and testing phase set to $384\times 128$ by default. To accelerate convergence, we adopt relatively large batch size (i.e., 128) and relatively large learning rate (i.e., 1e-4), which are different from original settings used in [19] (i.e., 64 and 1e-5).
Person attribute recognition. We use the PAR [40]as the baseline model, as implemented in [11]. The model is initialized with pretrained models and further fine-tuned on PA100 [40] and PETAzs [41] respectively. The image size is set to $256\times 128$, and all models are fine-tuned for 25 epochs by default.
Human pose estimation. We use the ViTPose [20] as the basic solution for HPE, implemented with open-source mmpose library [^2] https://github.com/open-mmlab/mmpose. The ViTPose model is initialized with pretrained models, fine-tuned on COCO keypoint [42] train2017, and evaluated on COCO keypoint val2017. In our experiments, the image size is set to $256\times 192$ and other settings follow the default configuration in mmpose.
Person landmarks detection. We use the mmpose to implement PLD model based on ViTPose, which is fine-tuned on whole-body COCO [43] train and evaluated on whole-body COCO val. The image size is set to $256\times 192$ in our experiments and other settings follow the default configuration in mmpose.
Body part segmentation. We implement body part segmentation model based on SCHP [22] method, and use the open-sourced code provided in [11] for training and testing. All pretrained models are fine-tuned on LIP [23] dataset and evaluated on LIP test set. Additionally, we adopt AdamW [45] as the optimizer during fine-tuning. Other implementation settings follow those of SCHP [22].
Pedestrian detection. We adopt CrowdDet [25] with pretrained backbone models as the basic solution for pedestrian detection. Those models are fine-tuned on CrowdHuman [44] using open-sourced mmdetection [^3] https://github.com/open-mmlab/mmdetection. We use the default configuration provided in mmdetection for training and testing, where AdamW optimizer with a learning rate of 2e-4 are used for training.
Multiple human parsing. We evaluate pretrained models on MHP using the Parsing-RCNN [21] implemented in [54]. The pretrained models are fine-tuned on CIHP [26] train for 25 training epochs (i.e., 1×1$\times$1 × schedule), and evaluated on CIHP val.
Part-level attribute parsing. We evaluate pretrained models on PAP using the KE-RCNN [24]. The pretrained models are fine-tuned on Fashionpedia [27] train for 32 training epochs (i.e., 1×1$\times$1 × schedule), and evaluated on Fashionpedia val.